# The GECKOS Survey: Identifying kinematic sub-structures in edge-on galaxies

### v1

@@
22institutetext: ARC Centre of Excellence for All Sky Astrophysics in Three Dimensions (ASTRO-3D) 33institutetext: School of Physics, University of New South Wales, NSW, 2052, Australia 44institutetext: Centre for Extragalactic Astronomy, Department of Physics, Durham University, South Road, Durham DH1 3LE, UK 55institutetext: National Research Council of Canada, Herzberg Astronomy and Astrophysics Research Centre, 5071 W. Saanich Rd. Victoria, BC, V9E 2E7, Canada 66institutetext: Centre for Astrophysics and Supercomputing, Swinburne University of Technology, PO Box 218, Hawthorn, VIC 3122, Australia 77institutetext: Astrophysics Research Institute, Liverpool John Moores University, 146 Brownlow Hill, Liverpool L3 5RF, UK 88institutetext: Sub-department of Astrophysics, Department of Physics, University of Oxford, Denys Wilkinson Building, Keble Road, Oxford OX1 3RH 99institutetext: Max-Planck-Institut für Extraterrestrische Physik, Gießenbachstraße 1, 85748 Garching, Germany 1010institutetext: International Centre for Radio Astronomy Research (ICRAR), The University of Western Australia, M468, 35 Stirling Highway, Crawley, WA 6009, Australia $1$institutetext: Research School of Astronomy and Astrophysics, Australian National University, Cotter Road, Weston Creek, ACT 2611, Australia 1212institutetext: Sydney Institute for Astronomy, School of Physics, A28, The University of Sydney, NSW, 2006, Australia 1313institutetext: Observatoire de Paris, LERMA, CNRS, PSL University, Sorbonne University, 75014 Paris, France 1414institutetext: Collège de France, 11 Pl. Marcelin Berthelot, 75231 Paris, France 1515institutetext: Cardiff Hub for Astrophysics Research & Technology, School of Physics & Astronomy, Cardiff University, Queens Buildings, Cardiff CF24 3AA, UK 1616institutetext: Instituto de Astrofísica de Canarias, Calle Vía Láctea s/n, E-38205 La Laguna, Tenerife, Spain 1717institutetext: Departamento de Astrofísica, Universidad de La Laguna, Av. del Astrofísico Francisco Sánchez s/n, E-38206, La Laguna, Tenerife, Spain 1818institutetext: Institute for Computational Cosmology, Department of Physics, Durham University, South Road, Durham DH1 3LE, UK 1919institutetext: Homer L. Dodge Department of Physics & Astronomy, University of Oklahoma, 440 W. Brooks St., Norman, OK 73019, USA 2020institutetext: Research Centre for Astronomy, Astrophysics, and Astrophotonics, Department of Physics and Astronomy, Macquarie University, NSW 2109, Australia 2121institutetext: European Southern Observatory, Alonso de Córdova 3107, Vitacura, Región Metropolitana $2$institutetext: Universidade Cidade de São Paulo/Universidade Cruzeiro do Sul, Rua Galvão Bueno 868, São Paulo-SP, 01506-000, Brazil 2323institutetext: Universitäts-Sternwarte, Fakultät für Physik, Ludwig-Maximilians-Universität München, Scheinerstr. 1, 81679 München, Germany 2424institutetext: Department of Astrophysics, University of Vienna, Türkenschanzstraße 17, 1180 Vienna, Austria 2525institutetext: Department of Physics and Astronomy, University of Utah, Salt Lake City, UT 84112, USA
@@
A. Fraser-McKelvie    J. van de Sande 1122    D. A. Gadotti 3322    E. Emsellem 44    T. Brown 11    D. B. Fisher 55    M. Martig 6622    M. Bureau 77    O. Gerhard 88    A. J. Battisti 99    J. Bland-Hawthorn 1010111122    B. Catinella 121222    F. Combes 101022    L. Cortese 13131414    S. M. Croom 101022    T. A. Davis 121222    J. Falcón-Barroso 1515    F. Fragkoudi 16161717    K. C. Freeman 1818    M. R. Hayden 111122    R. McDermid 191922    B. Mazzilli Ciraulo 202022    J. T. Mendel 5522    F. Pinna 111122    A. Poci 16161717    T. H. Rutherford 88    C. de Sá-Freitas 111212222121    L. A. Silva-Lima    L. M. Valenzuela $2$    G. van de Ven 2323    {CJK*}UTF8gbsn Z. Wang (王梓先) 2424    252522       A. B. Watts 101022
@@
The vertical evolution of galactic discs is governed by the sub-structures within them. Several of these features, including bulges and kinematically-distinct discs, are best studied in edge-on galaxies, as the viewing angle allows for easier separation of component light. In this work, we examine the diversity of kinematic sub-structure present in the first 12 galaxies observed from the GECKOS survey, a VLT/MUSE large programme providing a systematic study of 35 edge-on, Milky Way-mass disc galaxies. Employing the nGIST analysis pipeline, we derive the mean luminosity-weighted line-of-sight stellar velocity ($V_{\star}$), velocity dispersion ($\sigma_{\star}$), skew ($h_{3}$), and kurtosis ($h_{4}$) for the sample, and examine 2D maps and 1D line profiles. Scrutinising visual morphologies, the majority of this sample (8/12) are found to possess boxy-peanut bulges and host the corresponding kinematic structure predicted for stellar bars viewed in projection. Inferences are made on the bar viewing angle with respect to the line of sight from the strength of these kinematic indicators, finding one galaxy whose bar is close to side-on with respect to the observer, and two that are close to end-on. Four galaxies exhibit strong evidence for the presence of nuclear discs, including central $h_{3}$-$V_{\star}$ anti-correlations, ‘croissant’-shaped central depressions in $\sigma_{\star}$ maps, strong gradients in $h_{3}$, and positive $h_{4}$ plateaus over the expected nuclear disc extent. The strength of the $h_{3}$ feature corresponds to the size of the nuclear disc, measured from the $h_{3}$ turnover radius, taking into account geometric effects. We can explain the features within the kinematic maps of the four unbarred galaxies via disc structure(s) alone. We do not find any need to invoke the existence of dispersion-dominated bulges in any of the sample galaxies. Obtaining the specialised data products for this paper and the broader GECKOS survey required significant development of existing integral field spectroscopic (IFS) analysis tools. Therefore, we also present the nGIST pipeline: a modern, sophisticated, and easy-to-use pipeline for the analysis of galaxy IFS data, and the key tool employed by the GECKOS survey for producing value-added data products. We conclude that the variety of kinematic sub-structures seen in GECKOS galaxies requires a contemporary view of galaxy morphology, expanding on the traditional view of galaxy structure, and uniting the kinematic complexity observed in the Milky Way with the extragalactic.
@@ 1 Introduction
The Milky Way itself is largely considered to be a typical disc galaxy (e.g. Bland-Hawthorn & Gerhard, 2017). In addition to the structural components listed above, the Milky Way’s disc may be separated into two chemically distinct discs, whose stellar populations can be divided into young, metal-rich and [$\alpha$/Fe]-poor, and old, metal-poor and [$\alpha$/Fe]-rich (e.g. Gilmore & Reid, 1983; Haywood et al., 2013; Hayden et al., 2015). Observations of external galaxies have hinted at similar chemical structures (e.g. Pinna et al., 2019b, a; Scott et al., 2021). The kinematic thick disc of the Milky Way has also been found to host stars with a larger velocity dispersion and slower net rotation than stars in the thin disc (e.g. Chiba & Beers, 2000; Soubiran et al., 2003; Girard et al., 2006). Similar vertical dispersion gradients are also seen in external galaxies (e.g. Yoachim & Dalcanton, 2008; Comerón et al., 2019; Pinna et al., 2019b, a; Kasparova et al., 2020; Martig et al., 2021; Bhattacharya et al., 2023). Lastly, the Milky Way also hosts a bar, and a boxy-peanut (BP) bulge (e.g. Kent et al., 1991; Wegg & Gerhard, 2013; Ness & Lang, 2016). As such, the integrated kinematic measurements along a line of sight in an extragalactic Milky Way-like galaxy will include contributions from at least four different kinematic sub-structures near the centre.
@@
Significant progress in quantifying the motions of stars within BP bulges has been made, both in- and off-plane. Several studies have applied stellar kinematic indicators to classify the behaviour of the stars (e.g. Chung & Bureau, 2004; Bureau & Athanassoula, 2005; Méndez-Abreu et al., 2008, 2014), and dynamically, the bar orbits are well-understood and modelled. Barred galaxies contain stars on more elliptical orbits: chiefly the $x_{1}$ elongated orbits coincident with the canonical bar shape, and the $x_{2}$ orbits, which are smaller, generally more circular, and oriented roughly perpendicular to the $x_{1}$ orbits (e.g. Athanassoula, 1992). The angle at which the bar is viewed by the observer (i.e. more side-on or more end-on) determines both i) the relative contribution of each bar orbit family to a given line-of-sight velocity distribution (LOSVD) and ii) the shape of the LOSVD of that orbit (as bar orbits are not axisymmetric), hence the bar viewing angle has a measurable effect on the stellar kinematics observed.
@@
Some galactic features are difficult to model. The small scales of some central galactic features such as nuclear discs (Gadotti et al., 2020) have precluded them from being resolved in cosmological or even zoom simulations. Numerical heating of stellar particles also wreaks havoc on the size and shape of small, thin discs, even at high resolutions (e.g. Ludlow et al., 2019; Anta et al., 2023). Furthermore, most mock observations from simulations tend not to include typical observational effects such as dust lanes, although some advances have been made in this area (e.g. Barrientos Acevedo et al., 2023). Given the prevalence and obvious complexity of galactic kinematic sub-structures coupled with both simulation predictions and limitations, a census of kinematic sub-structures in a representative sample of Milky Way-mass galaxies will allow us to make progress towards a better interpretation and framework of all central disc structures.
In this paper, we categorise the diversity of kinematic sub-structures in the first 12 galaxies observed as part of the GECKOS [^2] Generalising Edge-on galaxies and their Chemical bimodalities, Kinematics, and Outflows out to Solar environments survey. We aim to define a range of kinematic sub-structures and link them to the assembly and evolutionary history of the host galaxies. In Section 2 we describe the GECKOS survey (including data reduction and analysis) and ancillary data. In Section 3 we identify BP bulges from imaging and present 1D and 2D stellar kinematics. In Section 4 we interpret the results within a framework of galactic central structure assembly. We also introduce the nGIST package in Appendix A: a modern, freely-available IFS data analysis package, flexible enough to analyse any galaxy IFS data. Throughout this paper, we use $\Lambda$CDM cosmology, with $\Omega_{\rm{m}}=0.30$, $\Omega_{\rm{\Lambda}}=0.70$, and $H_{0}=70$ km s-1 Mpc-1, and a Chabrier (2003) initial mass function.
@@ 2 Data
Figure 1: $g^{\prime}ri$ colour images of the sample galaxies reconstructed from the MUSE data cubes of the first 12 galaxies observed as part of the GECKOS survey. In each case, the mosaicked cube has been rotated and aligned such that the dust lane (and photometric major axis) is horizontal. The GECKOS pointing strategy was such that the entire (visible) bulge was always included within the MUSE field of view, and at least one side of the galactic disc. Here we show only regions where the signal-to-noise ratio is greater than 2.5; the MUSE field of view often extends further vertically.
Our sample consists of the first 12 galaxies from the GECKOS survey, an ESO/VLT large programme awarded 317 hours on MUSE (van de Sande et al., 2023). In Section 2.1-2.3 we summarise the sample selection, observing strategy, and data reduction respectively, but a full description will be presented in van de Sande et al. (in prep.). We also introduce the updated data analysis pipeline package nGIST in Section 2.4 and Appendix A.
@@ 2.1 GECKOS sample
GECKOS aims to investigate the relative importance of the internal and external physical processes that drive the evolution of disc galaxies. The GECKOS MUSE survey consists of 35 edge-on disc galaxies, targeting out to a surface brightness of $\mu_{V}=23.5\leavevmode\nobreak\ \rm{mag}\leavevmode\nobreak\ \rm{arcsec}^{-2}$, similar to the Sun’s position in the Galactic disc (Melchior et al., 2007). Targets were selected within a distance range of $7010<D\leavevmode\nobreak\ [\textrm{Mpc}]<$ 70, 24 of which were sourced from the S4G survey (Sheth et al., 2010) and the remaining 11 from HyperLeda (Makarov et al., 2014). Seven of these galaxies that meet our sample selection criteria possess MUSE archival observations.
Galaxies with a stellar mass within ±$\pm 0.3$ dex of the Milky Way ($5\times 10^{10}\rm{M}_{\odot}$; Bland-Hawthorn & Gerhard, 2017) that are edge-on (or highly inclined) are targeted. Furthermore, GECKOS aims to include a diversity of morpho-kinematic and star-formation properties by selecting galaxies within a 2 dex range of mid-IR-derived star-formation rates (SFRs; Cutri et al., 2013; Leroy et al., 2021). The GECKOS targets are evenly divided into three bins of SFR, comprising a sub-main sequence, a main sequence, and a super-main sequence sample. This spread in SFR not only maximises the variety of possible assembly histories, but also increases the probability of detecting outflows, present in $\sim 30-50\%$ of main sequence galaxies (Stuber et al., 2021) and ubiquitous in starbursts (e.g. Veilleux et al., 2020).

### v2

@@
22institutetext: ARC Centre of Excellence for All Sky Astrophysics in Three Dimensions (ASTRO-3D) 33institutetext: School of Physics, University of New South Wales, NSW, 2052, Australia 44institutetext: Centre for Extragalactic Astronomy, Department of Physics, Durham University, South Road, Durham DH1 3LE, UK 55institutetext: National Research Council of Canada, Herzberg Astronomy and Astrophysics Research Centre, 5071 W. Saanich Rd. Victoria, BC, V9E 2E7, Canada 66institutetext: Centre for Astrophysics and Supercomputing, Swinburne University of Technology, PO Box 218, Hawthorn, VIC 3122, Australia 77institutetext: Astrophysics Research Institute, Liverpool John Moores University, 146 Brownlow Hill, Liverpool L3 5RF, UK 88institutetext: Sub-department of Astrophysics, Department of Physics, University of Oxford, Denys Wilkinson Building, Keble Road, Oxford OX1 3RH 99institutetext: Max-Planck-Institut für Extraterrestrische Physik, Gießenbachstraße 1, 85748 Garching, Germany 1010institutetext: International Centre for Radio Astronomy Research (ICRAR), The University of Western Australia, M468, 35 Stirling Highway, Crawley, WA 6009, Australia $1$institutetext: Research School of Astronomy and Astrophysics, Australian National University, Cotter Road, Weston Creek, ACT 2611, Australia 1212institutetext: Sydney Institute for Astronomy, School of Physics, A28, The University of Sydney, NSW, 2006, Australia 1313institutetext: Department of Astrophysics, University of Vienna, Türkenschanzstraße 17, 1180 Vienna, Austria 1414institutetext: Observatoire de Paris, LERMA, CNRS, PSL University, Sorbonne University, 75014 Paris, France 1515institutetext: Collège de France, 11 Pl. Marcelin Berthelot, 75231 Paris, France 1616institutetext: Cardiff Hub for Astrophysics Research & Technology, School of Physics & Astronomy, Cardiff University, Queens Buildings, Cardiff CF24 3AA, UK 1717institutetext: Instituto de Astrofísica de Canarias, Calle Vía Láctea s/n, E-38205 La Laguna, Tenerife, Spain 1818institutetext: Departamento de Astrofísica, Universidad de La Laguna, Av. del Astrofísico Francisco Sánchez s/n, E-38206, La Laguna, Tenerife, Spain 1919institutetext: Institute for Computational Cosmology, Department of Physics, Durham University, South Road, Durham DH1 3LE, UK 2020institutetext: Homer L. Dodge Department of Physics & Astronomy, University of Oklahoma, 440 W. Brooks St., Norman, OK 73019, USA 2121institutetext: Research Centre for Astronomy, Astrophysics, and Astrophotonics, Department of Physics and Astronomy, Macquarie University, NSW 2109, Australia $2$institutetext: European Southern Observatory, Alonso de Córdova 3107, Vitacura, Región Metropolitana 2323institutetext: Universidade Cidade de São Paulo/Universidade Cruzeiro do Sul, Rua Galvão Bueno 868, São Paulo-SP, 01506-000, Brazil 2424institutetext: Universitäts-Sternwarte, Fakultät für Physik, Ludwig-Maximilians-Universität München, Scheinerstr. 1, 81679 München, Germany 2525institutetext: Department of Physics and Astronomy, University of Utah, Salt Lake City, UT 84112, USA
@@
A. Fraser-McKelvie    J. van de Sande 1122    D. A. Gadotti 3322    E. Emsellem 44    T. Brown 11    D. B. Fisher 55    M. Martig 6622    M. Bureau 77    O. Gerhard 88    A. J. Battisti 99    J. Bland-Hawthorn 1010111122    A. Boecker 121222    B. Catinella 1313    F. Combes 101022    L. Cortese 14141515    S. M. Croom 101022    T. A. Davis 121222    J. Falcón-Barroso 1616    F. Fragkoudi 17171818    K. C. Freeman 1919    M. R. Hayden 111122    R. McDermid 202022    B. Mazzilli Ciraulo 212122    J. T. Mendel 5522    F. Pinna 111122    A. Poci 17171818    T. H. Rutherford 88    C. de Sá-Freitas 111212222222    L. A. Silva-Lima    L. M. Valenzuela 2323    G. van de Ven 2424    {CJK*}UTF8gbsn Z. Wang (王梓先) 1313    252522       A. B. Watts 101022
@@
The vertical evolution of galactic discs is governed by the sub-structures within them. Several of these features, including bulges and kinematically-distinct discs, are best studied in edge-on galaxies, as the viewing angle allows for easier separation of component light. In this work, we examine the diversity of kinematic sub-structure present in the first 12 galaxies observed from the GECKOS survey, a VLT/MUSE large programme providing a systematic study of 36 edge-on, Milky Way-mass disc galaxies. Employing the nGIST analysis pipeline, we derive the mean luminosity-weighted line-of-sight stellar velocity ($V_{\star}$), velocity dispersion ($\sigma_{\star}$), skew ($h_{3}$), and kurtosis ($h_{4}$) for the sample, and examine 2D maps and 1D line profiles. Common clear kinematic signatures are observed: all galaxies display $h_{3}$ $V_{\star}$ sign mismatches in outer disc regions consistent with a (quasi-)axisymmetric, rotating disc of stars. Scrutinising visual morphologies, the majority of this sample (8/12) are found to possess boxy-peanut bulges and host the corresponding kinematic structure predicted for stellar bars viewed in projection. Inferences are made on the bar viewing angle with respect to the line of sight from the strength of these kinematic indicators, finding one galaxy whose bar is close to side-on with respect to the observer, and two that are close to end-on. Four galaxies exhibit strong evidence for the presence of nuclear discs, including central $h_{3}$-$V_{\star}$ profile anti-correlations, ‘croissant’-shaped central depressions in $\sigma_{\star}$ maps, strong gradients in $h_{3}$, and positive $h_{4}$ plateaus over the expected nuclear disc extent. The strength of the $h_{3}$ feature corresponds to the size of the nuclear disc, measured from the $h_{3}$ turnover radius, taking into account geometric effects. We can explain the features within the kinematic maps of the four unbarred galaxies via disc structure(s) alone. We do not find any need to invoke the existence of dispersion-dominated bulges in any of the sample galaxies. Obtaining the specialised data products for this paper and the broader GECKOS survey required significant development of existing integral field spectroscopic (IFS) analysis tools. Therefore, we also present the nGIST pipeline: a modern, sophisticated, and easy-to-use pipeline for the analysis of galaxy IFS data, and the key tool employed by the GECKOS survey for producing value-added data products. We conclude that the variety of kinematic sub-structures seen in GECKOS galaxies requires a contemporary view of galaxy morphology, expanding on the traditional view of galaxy structure, and uniting the kinematic complexity observed in the Milky Way with the extragalactic.
@@ 1 Introduction
The Milky Way itself is largely considered to be a typical disc galaxy (e.g. Bland-Hawthorn & Gerhard, 2016). In addition to the structural components listed above, the Milky Way’s disc may be separated into two chemically distinct discs, whose stellar populations can be divided into young, metal-rich and [$\alpha$/Fe]-poor, and old, metal-poor and [$\alpha$/Fe]-rich (e.g. Gilmore & Reid, 1983; Haywood et al., 2013; Hayden et al., 2015). Observations of external galaxies have hinted at similar chemical structures (e.g. Pinna et al., 2019b, a; Scott et al., 2021). The kinematic thick disc of the Milky Way has also been found to host stars with a larger velocity dispersion and slower net rotation than stars in the thin disc (e.g. Chiba & Beers, 2000; Soubiran et al., 2003; Girard et al., 2006). Similar vertical dispersion gradients are also seen in external galaxies (e.g. Yoachim & Dalcanton, 2008; Comerón et al., 2019; Pinna et al., 2019b, a; Kasparova et al., 2020; Martig et al., 2021; Bhattacharya et al., 2023). Lastly, the Milky Way also hosts a bar, and a boxy-peanut (BP) bulge (e.g. Kent et al., 1991; Wegg & Gerhard, 2013; Ness & Lang, 2016). As such, the integrated kinematic measurements along a line of sight in an extragalactic Milky Way-like galaxy will include contributions from at least four different kinematic sub-structures near the centre.
@@
Significant progress in quantifying the motions of stars within BP bulges has been made, both in- and off-plane. Several studies have applied stellar kinematic indicators to classify the behaviour of the stars (e.g. Chung & Bureau, 2004; Bureau & Athanassoula, 2005; Méndez-Abreu et al., 2008, 2014), and dynamically, the bar orbits are well-understood and modelled. Barred galaxies contain stars on more elongated orbits: chiefly the $x_{1}$ orbits coincident with the canonical bar shape, and the $x_{2}$ orbits, which are smaller, generally more circular, and oriented perpendicular to the $x_{1}$ orbits (e.g. Athanassoula, 1992; Sellwood & Wilkinson, 1993). The angle at which the bar is viewed by the observer (i.e. more side-on or more end-on) determines both i) the relative contribution of each bar orbit family to a given line-of-sight velocity distribution (LOSVD) and ii) the shape of the LOSVD of that orbit (as bar orbits are not axisymmetric), hence the bar viewing angle has a measurable effect on the stellar kinematics observed.
@@
Some galactic features are difficult to model. The small scales of some central galactic features such as nuclear discs (Gadotti et al., 2020) have precluded them from being resolved in cosmological or even zoom simulations. Numerical heating of stellar particles also wreaks havoc on the size and shape of small, thin discs, even at high resolutions (e.g. Ludlow et al., 2019; Anta et al., 2023). Furthermore, most mock observations from simulations tend not to include typical observational features such as dust lanes, although some advances have been made in this area (e.g. Barrientos Acevedo et al., 2023). Given the prevalence and obvious complexity of galactic kinematic sub-structures coupled with both simulation predictions and limitations, a census of kinematic sub-structures in a representative sample of Milky Way-mass galaxies will allow us to make progress towards a better interpretation and framework of all central disc structures.
In this paper, we categorise the diversity of kinematic sub-structures in the first 12 galaxies observed as part of the GECKOS [^2] Generalising Edge-on galaxies and their Chemical bimodalities, Kinematics, and Outflows out to Solar environments survey. We aim to define a range of kinematic sub-structures and link them to the assembly and evolutionary history of the host galaxies. In Section 2 we describe the GECKOS survey (including data reduction and analysis) and ancillary data. In Section 3 we identify BP bulges from imaging and present 1D and 2D stellar kinematics. In Section 4 we interpret the results within a framework of galactic central structure assembly. We also introduce the nGIST [^3] https://github.com/geckos-survey/ngist package in Appendix A: a modern, freely-available IFS data analysis package, flexible enough to analyse any galaxy IFS data. Throughout this paper, we use $\Lambda$CDM cosmology, with $\Omega_{\rm{m}}=0.30$, $\Omega_{\rm{\Lambda}}=0.70$, and $H_{0}=70$ km s-1 Mpc-1, and a Chabrier (2003) initial mass function.
@@ 2 Data
Figure 1: $g^{\prime}ri$ colour images of the sample galaxies reconstructed from the MUSE data cubes of the first 12 galaxies observed as part of the GECKOS survey. In each case, the mosaicked cube has been rotated and aligned such that the photometric major axis is horizontal, and if the dust lane is offset from it, it is always on the lower side of the image. For galaxies not exactly edge-on this means that we look at them from above (i.e. the lower half is the near half) if the dust lane is assumed to be in the midplane of the galaxy. The GECKOS pointing strategy was such that the entire (visible) bulge was always included within the MUSE field of view, and at least one side of the galactic disc. Here we show only regions where the signal-to-noise ratio is greater than 2.5; the MUSE field of view often extends further vertically.
Our sample consists of the first 12 galaxies from the GECKOS survey, an ESO/VLT large programme awarded 317 hours on MUSE (van de Sande et al., 2024). In Section 2.1-2.3 we summarise the sample selection, observing strategy, and data reduction respectively, but a full description will be presented in van de Sande et al. (in prep.). We also introduce the updated data analysis pipeline package nGIST in Section 2.4 and Appendix A.
@@ 2.1 GECKOS sample
GECKOS aims to investigate the relative importance of the internal and external physical processes that drive the evolution of disc galaxies. The GECKOS MUSE survey consists of 36 edge-on disc galaxies, covering out to a surface brightness of $\mu_{V}=23.5\leavevmode\nobreak\ \rm{mag}\leavevmode\nobreak\ \rm{arcsec}^{-2}$, similar to that at the Sun’s position in the Galactic disc when viewed face-on (Melchior et al., 2007). Targets were selected within a distance range of $7010<D\leavevmode\nobreak\ [\textrm{Mpc}]<$ 70, 25 of which were sourced from the S4G survey (Sheth et al., 2010) and the remaining 11 from HyperLeda (Makarov et al., 2014). Seven of these galaxies that meet our sample selection criteria possess MUSE archival observations.
Galaxies with a stellar mass within ±$\pm 0.3$ dex of the Milky Way ($5\times 10^{10}\rm{M}_{\odot}$; Bland-Hawthorn & Gerhard, 2016) that are edge-on (or highly inclined) are targeted. Furthermore, GECKOS aims to include a diversity of morpho-kinematic and star-formation properties by selecting galaxies within a 2 dex range of mid-IR-derived star-formation rates (SFRs; Cutri et al., 2013; Leroy et al., 2021). The GECKOS targets are evenly divided into three bins of SFR, comprising a sub-main sequence, a main sequence, and a super-main sequence sample. This spread in SFR not only maximises the variety of possible assembly histories, but also increases the probability of detecting outflows, present in $\sim 30-50\%$ of main sequence galaxies (Stuber et al., 2021) and ubiquitous in starbursts (e.g. Veilleux et al., 2020).
@@
In this paper, we use all galaxies whose observations were completed by December 2023, for either the entire galaxy or the central pointing. This subset of galaxies forms the core of the GECKOS first internal data release (iDR1) and includes fully reduced and mosaicked data cubes for the galaxies ESO 484-036, NGC 5775, ESO 079-003, UGC 00903, ESO 120-016, NGC 3279, NGC 0360, PGC 044931, IC 1711, and NGC 3957, and the central pointing only of NGC 0522 and IC 5244. While not a statistical nor representative sample of the local Universe, this sub-sample contains galaxies from all three bins of star formation rate, and is hence a representative sample of GECKOS galaxies.